OpenStudy (anonymous):

factor the trinomial 3u^2+1u-24

OpenStudy (id21):

First find the roots: \[u_{1,2}=-3; 8/3\] Then make a factor (x-x1)(x-x2) \[(x+3)(x-8/3)\]

OpenStudy (id21):

Do you know how to find the roots of equation \[3u^2+u-24=0\]?

OpenStudy (anonymous):

no

OpenStudy (id21):

\[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] For your case: \[3u^2+u-24=0\] \[u_{1,2}=\frac{-1 \pm \sqrt{1^2-4 \times 3 \times (-24)}}{2\times3}=\frac{-1\pm \sqrt{289}}{6}\] \[u_{1,2}=\frac{-18}{6}, \frac{16}{6}\]

OpenStudy (radar):

Suggest factoring: (u+3)(3u-8) set to 0 solve getting u=-3, and u=8/3

OpenStudy (radar):

same results as id21, but a different method.